OpenStudy (anonymous):

The length of a rectangular backyard is 13 feet longer than its width (x). What is the area of the backyard in the terms of x? x^2+13 13x^2+169x x^2+26x+169 x^2+13x

OpenStudy (anonymous):

13^2+169x

OpenStudy (anonymous):

How'd you get that answer?

OpenStudy (anonymous):

you don't want me to explain it it was confusing

OpenStudy (anonymous):

Nah, you just guessed.

OpenStudy (anonymous):

Oh well, we'll figure out if it was right or not.

OpenStudy (anonymous):

no i took 13 and squared it and then squared x because what you do to one side of the equation you have to do to the other and then i multiplied x to 169.

OpenStudy (anonymous):

The answer is the last option x^2+13x Let breadth=x Therefore Length=x+13 Area of a rectangle=Length*Breadth=(x)(x+13)=x^2+13x

OpenStudy (anonymous):

That's all you had to say! It's not confusing.

OpenStudy (anonymous):

Omg.

OpenStudy (anonymous):

well idk what im sayin so i leave

OpenStudy (anonymous):

@Kemist Hey, did you understand the solution?

OpenStudy (anonymous):

nope not really

OpenStudy (anonymous):

No, Sid. You'll have to explain later when I finish this test.

OpenStudy (anonymous):

Okay, let me know when you're done with it. Will surely help you with it.